OpenStudy (anonymous):

help solve: ln(2-ln3x) = 1

OpenStudy (tkhunny):

Transform! Use this \(log_{b}(a) = c \iff b^{c} = a\) for suitable a, b, and c.

OpenStudy (anonymous):

i did! and it gave me: \[2-\ln3x = e\] and then i put it to: \[2-e = \ln3x\] what now?! :O

OpenStudy (tkhunny):

What's preventing you from doing the same thing?

OpenStudy (anonymous):

so, x = 3 / e^2-e ?

OpenStudy (tkhunny):

Well, that's kind of horrible notation. Add some parentheses and you'll be closer. Get the division right-side up and you'll be golden. \(2−e=ln(3x) \implies e^{2-e} = 3x \implies x = \dfrac{e^{2-e}}{3}\)

OpenStudy (anonymous):

oh fail, I've been doing this assignment for 6 hours, I musn't have realized that I did a fail of a division there. thank you for pointing it out.

OpenStudy (tkhunny):

No worries. You were getting it. No trouble bailing you out after 6 lonely hours. Can't be there on your exam. Get enough rest and BE CAREFUL.

OpenStudy (anonymous):

thanks :)